# PromptDresser: Improving the Quality and Controllability of Virtual Try-On via Generative Textual Prompt and Prompt-aware Mask

### v1

@@ A Details of LMM-driven Virtual Try-on Captioning
We provide detailed explanations of the exemplar datasets, task descriptions, and templates for the categories of upper body, lower body, and dresses in Fig. 89, and 10, respectively. We first gave the LMM model the instruction to identify and list detailed attributes of a given person image, including components, such as the facial expression, skin color, clothing logos. We then selected attributes related to the masked regions of the person image or associated with the style of wearing the clothing, such as pose, hair length, and tucking style. For clothing images, we excluded attributes describing fine details, such as logo shapes or patterns, but instead focused on high-level attributes, such as the clothing category or sleeve.
@@ B Additional Details on User Study
Ours w/ LLaVA 0.8533 0.1222 8.96 1.04
Ours w/ GPT-4o 0.8530 0.116 8.64 0.75
@@
C Additional Experimental Results
@@
Additional Qualitative Comparisons. We present additional qualitative results in Fig. 11 and 12. The first three rows in Fig. 11 depict generated images on the VITON-HD [11] dataset using a model trained on the same dataset, while the fourth and fifth rows show generated images on the SHHQ-1.0 [16] dataset. Our model consistently generates the most realistic images, even for complex poses (rows 1 and 2), and addresses the issue of following the shape of the original clothing (rows 3 and 4). Notably, in the third row, only our model accurately captures the shape of the given cropped top. Moreover, Fig. 12 shows additional results for the upper body, lower body, and dresses categories on the DressCode [37] dataset. Similar to the results on the VITON-HD dataset, our  PromptDresser accurately generate the length of the clothing and mitigate the constraint the model follows the original clothing’s shape, highlighting the effectiveness of our rich text prompts and a novel mask refinement process.
Additional text-based editing Results. Fig. 13 and 14 demonstrate the text-editing capability of our PromptDresser on VITON-HD and lower body category of the DressCode datasets, respectively. Fig. 13 shows variations in tucking styles on the VITON-HD dataset, where the given clothing is generated based on the text prompts “fully tucked in”, “untucked”, and “french tuck”. Fig. 14 presents variations on the DressCode dataset, including “loose fit,” “tight fit,” and “pants rolled up.” The generated results demonstrate accurate and text-based editing capability of PromptDresser.

### v2

@@ 5 Conclusion
Acknowledgments This work was supported by Institute for Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2019-II190075, Artificial Intelligence Graduate School Program (KAIST)). This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2025-00555621). This work was supported by Technological Innovation R&D Program (RS-2024-00438252) funded by the Ministry of SMEs and Startups (MSS, Korea).
@@ A Details of LMM-driven Virtual Try-on Captioning
We provide detailed explanations of the exemplar datasets, task descriptions, and templates for the categories of upper body, lower body, and dresses in Fig. 1112, and 13, respectively. We first gave the LMM model the instruction to identify and list detailed attributes of a given person image, including components, such as the facial expression, skin color, clothing logos. We then selected attributes related to the masked regions of the person image or associated with the style of wearing the clothing, such as pose, hair length, and tucking style. For clothing images, we excluded attributes describing fine details, such as logo shapes or patterns, but instead focused on high-level attributes, such as the clothing category or sleeve.
@@ B Additional Details on User Study
Ours w/ LLaVA 0.8663 0.1175 8.85 0.91
Ours w/ GPT-4o 0.8686 0.1119 8.54 0.67
@@
C Experimental Details
[…]
D Additional Experimental Results
@@
Method sec/image SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ KID $\downarrow$
IDM-VTON (40step) 5.84 0.8613 0.1018 9.14 1.18
Ourspose 5.78 0.8778 0.0967 9.07 1.16
Ours 5.78 0.8686 0.1119 8.54 0.67
Table 7: Comparison of IDM-VTON at similar inference time.
Comparison of Inference Time. PMG derives a refined mask from the coarse mask during the initial denoising steps. While this process may introduce additional computational overhead, Table 7 shows that, when compared to IDM-VTON with slightly increased steps, our method demonstrates superior performance in the unpaired setting while maintaining comparable inference time.
GPT-Human Human-Human
STS 0.8622 0.8889
Table 8: STS between GPT-Human and Human-Human pairs.
Bias in LMM Evaluation. Recent studies [55, 20] validated LMMs for evaluation tasks such as image-to-text and multi-image-to-text alignment. Similarly, we used an LMM to evaluate outfit labeling. To verify the LMM-based evaluation, a user study (Fig. 8(c) in the main paper) assessed text-image alignment via human feedback, showing significant improvement over baselines. To verify LMM reliability, four human annotators labeled 100 VITON-HD test images. Table 8 shows mean semantic textual similarity (STS) between GPT and human labels (GPT-Human), as well as between all human-labeled pairs (Human-Human). Comparable GPT-Human and Human-Human scores confirm GPT’s alignment with human judgment.
Figure 9: Multi-layer / transparent outfit generation images.
SSIM $\downarrow$ LPIPS $\uparrow$
IDM-VTON 0.9401 0.0405
Ours 0.8702 0.1030
Table 9: Comparison of ‘tucked in’ vs. ‘untucked’.
Additional Visual Results and Diversity. Fig. 9 demonstrates outfit changes and transparent cases generated under various textual conditions, combining complex scenarios such as tops, bottoms, and outerwear. Although trained solely on the DressCode dataset, our method effectively edits images with multi-layered clothing and diverse outfits. We plan to update the image combinations in future versions. Table 9 compares SSIM and LPIPS between tucked and untucked generations. By leveraging rich text prompts and flexible mask, our method achieves higher editability than IDM-VTON.
Figure 10: Generation results of VITON-HD.
Additional Visual Results of Mask Augmentation. Fig. 10 shows the effect of mask augmentation. As noted in the paper, training solely with fine masks causes the model to fit the clothing strictly within the mask region. Without augmentation, it fails to handle coarse masks (e.g., the large rectangular mask), generating overly long garments, as shown in Fig. 10.
Additional Qualitative Comparisons. We present additional qualitative results in Fig. 14 and 15. The first three rows in Fig. 14 depict generated images on the VITON-HD [10] dataset using a model trained on the same dataset, while the fourth and fifth rows show generated images on the SHHQ-1.0 [15] dataset. Our model consistently generates the most realistic images, even for complex poses (rows 1 and 2), and addresses the issue of following the shape of the original clothing (rows 3 and 4). Notably, in the third row, only our model accurately captures the shape of the given cropped top. Moreover, Fig. 15 shows additional results for the upper body, lower body, and dresses categories on the DressCode [37] dataset. Similar to the results on the VITON-HD dataset, our  PromptDresser accurately generate the length of the clothing and mitigate the constraint the model follows the original clothing’s shape, highlighting the effectiveness of our rich text prompts and a novel mask refinement process.
Additional text-based editing Results. Fig. 16 and 17 demonstrate the text-editing capability of our PromptDresser on VITON-HD and lower body category of the DressCode datasets, respectively. Fig. 16 shows variations in tucking styles on the VITON-HD dataset, where the given clothing is generated based on the text prompts “fully tucked in”, “untucked”, and “french tucked”. Fig. 17 presents variations on the DressCode dataset, including “loose fit,” “tight fit,” and “pants rolled up.” The generated results demonstrate accurate and text-based editing capability of PromptDresser.